OpenStudy (anonymous):

Find the length and width of a rectangle that has an area of 512 square feet and whose perimeter is a minimum.

OpenStudy (anonymous):

so far i have xy=512 and 2x+2y=P, so 2(512/y)+2y=P

OpenStudy (anonymous):

Do you have multiple choice answers? Also it's not really possible to find it without a perimeter.

OpenStudy (anonymous):

no its not multiple choice. its optimization

OpenStudy (inkyvoyd):

Chuk, let me help.

OpenStudy (anonymous):

i think i'm supposed to take the derivative but i always get confused when the variable is in the denominator

OpenStudy (inkyvoyd):

2(512/y)+2y=P -> multiply everything by y. -> the result is a quadratic (P is a constan)

OpenStudy (inkyvoyd):

@ChukRock , you don't need derivatives, although your prof probably wants them.

OpenStudy (anonymous):

so i should get 1024+2y^2=p

OpenStudy (inkyvoyd):

btw, with the resulting quadratic, you'll want to set the derivative to 0, and do dP/dy, because P is a function f y.

OpenStudy (inkyvoyd):

*=py

OpenStudy (inkyvoyd):

Actually, that makes it messier. Don't multiply, sorry.

OpenStudy (anonymous):

i think taking the deriv got me (2048/y^2)+2=p

OpenStudy (inkyvoyd):

Can we redo the problem, and just get rid of the y? I'll teach you how to set it up.

OpenStudy (anonymous):

well i kinda want to learn the way of taking the derivative because i'm sure the prof will want to see work

OpenStudy (inkyvoyd):

I know, I mean resetting up the problem so you only need 2 variables.

OpenStudy (anonymous):

well at the point i got it to wouldnt i replace the p with zero and its only one variable?

OpenStudy (inkyvoyd):

Well, you actally did the your setup wrong. should've been py, not p.

OpenStudy (inkyvoyd):

Let me show you what I would do.

OpenStudy (anonymous):

ok

OpenStudy (inkyvoyd):

x(p/2-x)=512 Minimize P.

OpenStudy (anonymous):

i've never seen this way

OpenStudy (inkyvoyd):

Because, I got rid of the y.

OpenStudy (inkyvoyd):

P=2x+2y ->P/2=x+y -> P/2-x=y xy=512

OpenStudy (inkyvoyd):

The right substitution makes things easier. :)

OpenStudy (anonymous):

so (xp/2)-x^2=512? maybe it makes it easier for you but now i'm even more lost

OpenStudy (inkyvoyd):

find dP/dx and set it to zero. Use implicit differention as well as the product rule. If you haven't learned those, wel, I'll think of another way.

OpenStudy (anonymous):

the way i learned was to substitute either x=(512/y) or y=(512/x) into the objective function

OpenStudy (anonymous):

i'm just trying to figure out how to do it THAT way

OpenStudy (inkyvoyd):

That's all good, but getting that 1/y is harder to solve when you find the derivative.

OpenStudy (anonymous):

thank you anyway

OpenStudy (inkyvoyd):

in any case, you'd wind up with implicity differentiation. I just realized that your method is actually faster.

OpenStudy (inkyvoyd):

But, that doesn't keep you from having to use the product rule and implicit differentiaion.

OpenStudy (inkyvoyd):

the answer (obviously) is 16sqrt2

OpenStudy (anonymous):

@inkyvoyd i figured it out the way to do it with derrivs and you dont need chain or implicit diff\[2(512/y)+2y=P\]\[(1024/y)+2y=P\]\[-1024/y ^{2}+2=P\] at this point P becomes zero\[2=1024/y ^{2}\]\[2y ^{2}=1024\]\[y ^{2}=512\]\[y=16\sqrt{2}\]

OpenStudy (inkyvoyd):

That's half correct. You forgot to divide by y on the right hand side of the equation, but the answer still comes out.